Find an equation of the line containing the given air of points.
(-1,-3) and (-6,-5)

Answer by stanbon(75887)

Find an equation of the line containing the given pair of points.
(-1,-3) and (-6,-5)
-----------
slope = (-5--3)/(-6--1) = -2/-5 = 2/5
----
Equation Form:
y = mx + b
-3 = (2/5)(-1)+b
-15/5 = -2/5 + b
b = -13/5
========
Equation:
y = (2/5)x - (13/5)
=====================
Cheers,
Stan H.
=====================